# LLMs for Law: Evaluating Legal-Specific LLMs on Contract Understanding

### v1

@@ 5.8M T5-base
2.2 Tasks and Datasets
Table 2 summarizes key details. Appendix A provides statistics and examples of the datasets which are as follows:
UNFAIR-ToS The UNFAIR-ToS dataset from Chalkidis et al. (2022) is used to identify unfair contractual terms in Terms of Service (ToS) documents from platforms like YouTube, eBay, and Facebook. Each sentence is annotated with one or more of 8 unfairness categories, plus 1 unlabeled class for sentences that do not indicate any potential violation of European consumer law. This makes the task a multi-label classification problem. Labels are based on potential violations of EU consumer protection law . The dataset includes training (5.5k), development (2.3k), and test (1.6k) sets.
LEDGAR The LEDGAR dataset from Chalkidis et al. (2022) is used to classify the principal topic of provisions in Exhibit 10 material contracts (e.g., employment, lease, non-disclosure) filed with the US Securities and Exchange Commission (SEC) via EDGAR . Each provision (paragraph) is labeled with one of 100 contract topics, making it a multi-class classification task. The dataset includes training (60k), development (10k), and test (10k) sets.
LEXDEMOD The LEXDEMOD dataset from Sancheti et al. (2022) detects deontic modality in agent-based contract clauses from lease agreements sourced from the LEDGAR dataset. Each clause (sentence) is annotated with one or more of 6 deontic modality types plus 1 none class , making it a multi-label classification task. Labels are linked to an agent (party) in the sentence, representing their deontic status (e.g., Obligation, Entitlement, Prohibition). The dataset includes training (4.2k), development (330), and test (1.7k) sets. The train/validation/test split is as reported in the original paper.
3 Experiment Setup
We perform task-specific (supervised) fine-tuning using 10 legal-specific LLMs on three datasets: LEDGAR, UNFAIR-ToS, and LEXDEMOD. We consider 10 pre-trained encoder-based legal-specific models for fine-tuning. Nine of these are base-variant encoder models: Legal-BERT Chalkidis et al. (2020), Contracts-BERT Chalkidis et al. (2020), LegalRoBERTa Geng et al. (2021), CaseLaw-BERT Zheng et al. (2021), InLegalBERT, InCaseLawBERT, and CustomInLawBERT Paul et al. (2023), Legal-XLM-R Niklaus et al. (2024), and LexLM Chalkidis* et al. (2023). One large-variant model, PoL-BERT Henderson et al. (2022), is included, as its base version is not present. We also evaluate the encoder-decoder model LexT5 T.y.s.s et al. (2024) (Appendix E), but exclude it from the main results as it is the only model of its kind. Decoder-only models like AdaptLLM Cheng et al. (2024) and SaulLM-7B Colombo et al. (2024) are emerging but custom metrics are not well-supported by the TRL library, which we require in the contract classification case. We leave their benchmarking for future work. A detailed description of each model is provided in Appendix B, and Table 3 summarizes their key characteristics. A detailed experimental setup is provided in Appendix C.
We also compare the 10 legal-specific LLMs with 7 general-purpose LLMs. These include five base variant encoder models: BERT Devlin et al. (2019), RoBERTa-base Liu et al. (2019), DeBERTa He et al. (2021), Longformer Beltagy et al. (2020), and BigBird Zaheer et al. (2020), along with one large variant, RoBERTa-large Liu et al. (2019). Additionally, we compare with the closed-source GPT-3.5-Turbo OpenAI (2022) using zero-shot and one-shot prompting.
UNFAIR-ToS LEDGAR LEXDEMOD
Method Model # Params $\mu$-F1 m-F1 $\mu$-F1 m-F1 $\mu$-F1 m-F1
Zero-shot GPT-3.5-Turbo In Billions 41.4 22.2 70.1 56.7 - -
Few-shot GPT-3.5-Turbo In Billions 64.7 32.5 62.1 51.1 - -
BERT 110M 95.6 81.3 87.6 81.8 - 75.61
RoBERTa-base 125M 95.2 79.2 87.9 82.3 - 75.66
DeBERTa 139M 95.5 80.3 88.2 83.1 - -
Longformer 149M 95.5 80.9 88.2 83.0 - -
BigBird 127M 95.7 81.3 87.8 82.6 - -
Baselines reported
from: Chalkidis (2023),
Chalkidis et al. (2022),
Sancheti et al. (2022)
SFT
(General-
purpose
LLMs)
RoBERTa-large 355M 95.8 81.6 88.6 83.6 - 77.88
Legal-BERT 110M 96.0 82.2 88.2 82.5 81.23 78.01
Contracts-BERT 110M 96.2 83.4 87.9 82.2 80.17 77.71
Legal-RoBERTa 125M 95.4 81.1 87.7 81.9 80.12 76.70
CaseLawBERT 110M 96.1 83.2 87.6 80.9 80.32 77.75
PoL-BERT 340M 94.6 77.9 86.0 79.1 41.35 15.75
InLegalBERT 110M 95.6 81.7 87.9 82.0 80.21 77.89
InCaseLawBERT 110M 95.5 81.1 87.5 82.1 79.16 76.83
CustomInLawBERT 110M 95.5 79.9 87.7 81.8 78.16 75.35
LexLM 124M 95.9 81.7 87.8 81.3 80.39 77.46
Proposed
SFT
(Legal-
specific
LLMs)
Legal-XLM-R 184M 94.9 78.2 87.7 81.7 80.62 77.56
Table 4: Performance of legal-specific and general-purpose LLMs on three tasks: UNFAIR-ToS, LEDGAR, LEXDEMOD. Metrics: micro-F1 ($\mu$-F1) and macro-F1 (m-F1). SFT denotes supervised fine-tuning; zero-shot and few-shot indicate prompting methods. Red highlights best legal-specific, blue highlights best general-purpose performance.
Mean $\pm$ Std
Legal Specific
LLMs
$\mu$-F1 m-F1
Legal-BERT 88.48 ± 6.03 80.90 ± 2.05
Contracts-BERT 88.09 ± 6.55 81.10 ± 2.45
Legal-RoBERTa 87.74 ± 6.24 79.90 ± 2.29
CaseLawBERT 88.01 ± 6.45 80.62 ± 2.23
PoL-BERT 73.98 ± 23.34 57.58 ± 29.58
InLegalBERT 87.90 ± 6.28 80.53 ± 1.87
InCaseLawBERT 87.39 ± 6.67 80.01 ± 2.29
CustomInLawBERT 87.12 ± 7.09 79.02 ± 2.71
LexLM 88.03 ± 6.33 80.15 ± 1.91
Legal-XLM-R 87.74 ± 5.83 79.15 ± 1.82
LexT5 85.60 ± 7.73 76.40 ± 2.66
Table 5: Aggregated scores (Mean $\pm$ Std) across three contract understanding tasks. Red, blue, and green highlights indicate the first, second, and third best performances, respectively.
4 Results and Analysis
Table 4 reports the test results of LLMs across all three tasks, while Table 5 presents the aggregated scores. To address the main research question in Section 1, we run experiments to answer the following Sub-Research Questions (SRQs):
SRQ1: How do legal-specific LLMs perform across different contract understanding tasks compared to general-purpose LLMs?
Table 4 compares legal-specific and general-purpose LLMs. Among general models, RoBERTa-large (355M) performs best overall. However, LLMs such as Contracts-BERT and Legal-BERT (110M) outperform RoBERTa-large on UNFAIR-ToS and LEXDEMOD, respectively, despite having 69% fewer parameters. Other legal LLMs, including CaseLaw-BERT and LexLM, also surpass RoBERTa-large on UNFAIR-ToS. Legal-RoBERTa, CustomInLawBERT, InCaseLawBERT, and InLegalBERT consistently outperform RoBERTa-base, BERT, DeBERTa, and Longformer on UNFAIR-ToS. On LEXDEMOD, Legal-BERT and InLegalBERT again outperform RoBERTa-large. These results highlight that legal-specific base variant LLMs, despite having 64-69% fewer parameters, often outperform larger general-purpose LLMs on domain-specific tasks. RoBERTa-large remains the best model for LEDGAR. Still, Legal-BERT delivers equivalent performance compared to general-purpose base variant models on this task, suggesting that both model size and task characteristics influence performance. The larger legal-specific LLMs may be better suited for LEDGAR. Overall, we conclude that legal-specific base models deliver competitive performance and set new SOTAs on two of the three tasks, demonstrating the effectiveness of domain-specific pretraining, even at the base variant of LLMs.
SRQ2: Which legal-specific LLMs serve as strong baselines for contract understanding?
Table 5 presents aggregated test scores (arithmetic, harmonic, and geometric means) across the three contract understanding tasks. The top three performances are highlighted in red, blue, and green respectively. Despite class imbalance in all tasks, Legal-BERT achieves the highest aggregated $\mu$-F1, while Contracts-BERT leads in m-F1. Across both metrics, the top positions are consistently held by four legal-specific models: Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM. We conclude that these four models, Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM, should be considered strong baselines for contract understanding tasks.
SRQ3: What are the observed limitations of current legal-specific LLMs, and how can these findings guide future legal LLMs development?
Several recent legal-specific LLMs, such as Legal-RoBERTa, CaseLaw-BERT, PoL-BERT, CustomInLawBERT, LexLM, and Legal-XLM-R, are pre-trained on large-scale legal corpora. Models like InLegalBERT and InCaseLawBERT are built on legal-specific base models rather than general-purpose models. However, older legal-specific base-variant LLMs, such as LegalBERT and ContractsBERT, which are pre-trained on just 354k and 76k legal documents respectively (as seen in Table 3), still outperform many recent base-variant legal-specific models (as seen in Table 5). A key limitation of recent legal-specific LLMs is that they are pre-trained on few, or no, diverse contract documents compared to other legal texts like legislation and court cases. We conclude that future legal-specific LLMs should incorporate a more diverse and representative set of contract documents, to improve performance across contract understanding tasks.
5 Conclusion
This study benchmarks 10 legal-specific LLMs against 7 general-purpose LLMs across three contract understanding tasks. Legal-specific base LLMs consistently perform well and set new SOTAs on two tasks despite having fewer parameters. Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM emerge as strong baseline models for contract understanding. However, recent base-variant legal LLMs often underperform due to limited pretraining on diverse contract data. Future work focuses on expanding contract data and evaluating emerging decoder-based legal LLMs.
Limitations
The limited availability of contract benchmark datasets in languages other than English poses a challenge for multilingual extension. Consequently, this study focuses solely on English-language contract tasks, leaving evaluation on non-English data for future work. While encoder-decoder models like LexT5 and decoder-based legal-specific LLMs such as AdaptLLM and SaulLM-7B are emerging, they remain scarce. We therefore defer their benchmarking until more models become available, ensuring fair comparisons. LexT5 is evaluated for exploratory purposes but excluded from the main results. Additionally, this work concentrates on the nuances of contract language and does not assess performance on other legal text types, such as statutes, court decisions, or legal opinions. Future research should extend this evaluation to a broader range of legal genres, acknowledging that no single study can fully capture the entire legal domain.
Ethical Considerations
This study uses only publicly available datasets, LEDGAR, UNFAIR-ToS, and LEXDEMOD, all of which contain contract clauses without personal data. LEDGAR is derived from public U.S. SEC EDGAR filings, UNFAIR-ToS from company Terms of Service, and LEXDEMOD from lease clauses sourced from LEDGAR. This research does not offer legal advice, predict individual outcomes, or automate decisions affecting rights. It focuses solely on evaluating the performance of legal-specific LLMs to inform future tools and research. While these models can support legal professionals, they are not substitutes for legal expertise. We acknowledge potential ethical risks if outputs are misused or inaccurate. By open-sourcing our evaluations, we aim to reduce reliance on proprietary tools, promote transparency, and expand access to legal AI research and development.
Acknowledgment
We utilized commercial AI tools to enhance the clarity, grammar, and style of our manuscript.
Appendix A Dataset Statistics and Illustrative Examples
This appendix presents labeled examples from all three datasets to aid understanding. We provide our own rationales explaining label types, which the datasets do not explicitly include. These explanations clarify why specific labels apply to given clauses (sentences) or provisions (paragraphs), as detailed in Table 6. Legal contract classification involves longer texts than typical NLP tasks like tweets or reviews. Legal-specific Transformer models such as Legal-BERT process up to 512 sub-word tokens, but many LEDGAR paragraphs exceed this limit. Figure 1 shows that numerous LEDGAR paragraphs surpass the standard context window, requiring truncation or other methods to handle long inputs. Additionally, legal texts contain specialized terminology (legalese), increasing classification complexity compared to general text.
Dataset Example Label
Rationale for Assigned Labels
(Provided by us for better understanding)
This Amendment may be executed by one or more of the
parties hereto on any number of separate counterparts,
and all of said counterparts taken together shall be deemed
to constitute one and the same instrument. This
Amendment may be delivered by facsimile or other
electronic transmission of the relevant signature pages
hereof.
Counterparts
This sentence states that the Amendment may be
executed in multiple counterparts and that together
they form a single agreement, which is a standard
counterparts clause used to validate separately
signed copies as one binding document.
THIS AMENDMENT SHALL BE GOVERNED BY, AND
INTERPRETED IN ACCORDANCE WITH, THE LAW
OF THE STATE OF NEW YORK . The other provisions of
Article IX of the Credit Agreement shall apply to this
Amendment to the same extent as if fully set forth herein.
Governing Laws
This sentence specifies that New York law will
govern and interpret the Amendment, which is a
standard governing law clause that establishes
the legal jurisdiction and framework for resolving
disputes.
LEDGAR
Sublessee leases the Aircraft in its “as is, where is” condition.
The only services, rights, or warranties to which the Sublessee
is entitled to under this Sublease are those to which the
Sublessor is provided under the Prime Lease.
Warranties
The sentence is labeled as warranties because it
defines the rights and guarantees the Sublessee
receives and limits those warranties to what the
Sublessor has under the Prime Lease.
Niantic further reserves the right to remove any User
Content from the Service at any time and without notice
and for any reason.
Content removal
This sentence is labeled as content removal unfair
contractual term because it gives the provider full
control to remove content at any time, for any
reason, and without notice.
amazon reserves the right to refuse service, terminate
accounts , terminate your rights to use amazon services,
remove or edit content , or cancel orders in its sole
discretion.
Unilateral
termination,
Content removal
This sentence is labeled as unilateral termination and
content removal because it allows Amazon to end
services, remove content, or cancel orders at its sole
discretion, without notice, creating an imbalance of
power.
UNFAIR-ToS
Outside the United States and Canada. If you acquired
the if you acquired the application in any other country,
the laws of that country apply.
None
The sentence is labeled as none because it does not
belong to any of the unfair contractual term types
and is actually a fair clause.
[lessee] Lessee will not create or permit to be
created or to remain , and will promptly
discharge , any lien , encumbrance or charge (including
without limitation any mechanic ’s , laborer ’s or
materialman ’s lien ) against the Premises or any part
thereof arising from Lessee ’s actions.
Prohibition,
Obligation
This sentence imposes a prohibition by forbidding
the Lessee (the agent) from creating or allowing liens.
It also imposes an obligation by requiring the Lessee
to promptly remove any such liens.
[tenant] Tenant may, without Landlord’s consent,
before delinquency occurs, contest any such taxes related
to the Personal Property.
Permission
This sentence grants permission to the Tenant (the
agent) to contest taxes without needing the
Landlord’s consent, as long as it’s done before
delinquency.
LEXDEMOD
[landlord] Tenant shall promptly notify Landlord of any
alleged defaults under the CC&Rs and/or the Oil and Gas
Lease .
Entitlement
The Landlord, as the agent, holds an entitlement to
receive notice from the Tenant about alleged defaults.
Appendix B Description of Legal-specific LLMs
@@
UNFAIR-ToS LEDGAR LEXDEMOD
Method Model # Params $\mu$-F1 m-F1 $\mu$-F1 m-F1 $\mu$-F1 m-F1
Legal-BERT 110M 95.20 78.88 88.32 82.10 76.08 72.76
Contracts-BERT 110M 95.13 76.31 87.83 81.97 77.51 72.49
Legal-RoBERTa 125M 94.79 76.16 87.72 81.17 77.01 72.68
CaseLawBERT 110M 95.25 75.24 87.65 81.22 76.71 74.74
PoL-BERT 340M 93.92 68.80 85.76 78.60 48.65 17.43
InLegalBERT 110M 95.43 76.82 87.99 81.54 74.79 71.29
InCaseLawBERT 110M 95.44 75.87 87.41 80.52 78.01 76.75
CustomInLawBERT 110M 95.04 72.44 87.37 80.38 78.26 75.03
LexLM 124M 95.34 77.64 87.84 81.26 78.46 75.61
Proposed
SFT
(Legal-
specific
LLMs)
Legal-XLM-R 184M 94.90 71.89 87.50 81.02 77.59 73.50
Table 7: Validation results for all examined legal-specific LLMs. Model performance is evaluated using micro-F1 ($\mu$-F1) and macro-F1 (m-F1). SFT denotes Supervised Fine-Tuning.
@@ Appendix C Experimental Setup
We use all publicly available legal-specific pre-trained models from Hugging Face. To ensure fair comparison, we adopt the training configuration introduced by Chalkidis et al. (2022) for the LEDGAR and UNFAIR-ToS datasets: a learning rate of $3\text{e-}5$ for all nine encoder-base models and $1\text{e-}5$ for the encoder-large model PoL-BERT Henderson et al. (2022), consistent with the setting used for RoBERTa-large. All models are trained for up to 20 epochs with a batch size of 8, using early stopping with a patience of 3 based on development set performance. For UNFAIR-ToS, we use a maximum sequence length of 128, as in Chalkidis et al. (2022). However, we disable mixed-precision training (i.e., set fp16=False) to ensure stable training, which results in longer training times compared to Chalkidis et al. (2022). For LEDGAR, we reduce the maximum sequence length to 128 (from 512 used in Chalkidis et al. (2022)) to save computational resources and training time. This adjustment is necessary because, as discussed above, fp16 is disabled to ensure stable training, which leads to longer training times, and the LEDGAR dataset contains over 80k sentences, which is large. We observe only a marginal performance drop (0.1-0.4%), which we consider acceptable for efficiency. Each model is trained five times with different random seeds (1-5), and we report the test results of the best seed, following Chalkidis et al. (2022) for a fair baseline comparison. For the LEXDEMOD dataset, we follow the setup proposed by Sancheti et al. (2022), using a learning rate of $2\text{e-}5$ for all encoder-based legal-specific models, including PoL-BERT, consistent with their configuration for RoBERTa-large. We use a batch size of 8 and apply early stopping with a patience of 3. The maximum sequence length is set to 256, as in Sancheti et al. (2022). Each model is trained five times with different random seeds, and we report the average test performance across the three best seeds, following Sancheti et al. (2022) for a fair baseline comparison. We evaluate model performance using micro-F1 ($\mu$-F1) and macro-F1 (m-F1) to account for class imbalance. Additionally, we report the arithmetic mean with standard deviation for micro-F1 ($\mu$-F1) and macro-F1 (m-F1) across tasks. All experiments are conducted on a single NVIDIA V100 GPU. Although the datasets are already publicly available, we will release our code and evaluation scripts to support full reproducibility.
Appendix D Additional Result
Table 7 presents the development set results for all examined models across the three datasets. Validation results are reported in the same format as the test set, as detailed in the experimental setup in Appendix C.
Appendix E Other Legal-Specific LLMs Considered
In addition to encoder-based legal-specific LLMs, we experiment with encoder-decoder models, specifically LexT5, for exploratory purposes. Instead of supervised fine-tuning, we adopt instruction-based fine-tuning, which aligns better with encoder-decoder and decoder-only architectures. This approach pairs natural language prompts with clause inputs, enabling the model to generate the appropriate label(s) as output. The instruction templates used are listed in Figure 2. For UNFAIR-ToS, we directly use the prompt from Chalkidis (2023), originally designed for zero-shot prompting. For LEXDEMOD and LEDGAR, we design our own prompts inspired by that style. For our experiments, we use the LexT5 Base model with 220M parameters and apply the same hyperparameter configuration used for encoder-based models, as detailed in Appendix C, to ensure fair comparison. On the UNFAIR-ToS dataset, LexT5 achieves a micro-F1 of 95.4 and a macro-F1 of 79.8. While it does not outperform the encoder-based legal-specific model Contract-BERT, it performs better than the general-purpose baseline RoBERTa-large and surpasses some legal-specific models such as PoL-BERT and Legal-XLM-R. On the LEDGAR dataset, LexT5 achieves a micro-F1 of 84.9 and a macro-F1 of 76.1, which falls short of both general-purpose and legal-specific LLMs. On LEXDEMOD, it attains a micro-F1 of 76.5 and a macro-F1 of 73.3, failing to surpass any general-purpose LLMs and underperforming compared to the majority of encoder-based legal-specific models.
Overall, encoder-decoder legal-specific models such as LexT5 currently underperform compared to encoder-based legal-specific and general-purpose LLMs. However, drawing broad conclusions from a single encoder-decoder model is not fair. While encoder-decoder legal-specific models are beginning to emerge, they remain limited in number. At present, LexT5 T.y.s.s et al. (2024) is the only publicly available model of its kind that we are able to find. Decoder-based legal LLMs, such as AdaptLLM Cheng et al. (2024) and SaulLM 7B Colombo et al. (2024), are introduced in 2024; however, such models remain scarce. As a result, we defer comprehensive benchmarking of encoder-decoder and decoder-only legal-specific LLMs to future work, when a broader range of models becomes available to ensure fair comparison.

### v2

@@ 5.8M T5-base
AdaptLLM
Cheng et al. (2024)
US court opinions from FreeLaw project
3.6M LLama-1
SaulLM-7B
Cheng et al. (2024)
FreeLaw, English MultiLegal Pile, EDGAR, EuroParl, GovInfo
Law Stack Exchange, Australian legal corpora, EU and UK
legislation, court transcripts, and USPTO filings
- Mistral
@@ 3.6M LLama-1
Table 5 summarizes their key characteristic. The detailed description of the legal-specific models used in the experiments is given below:
@@
AdaptLLM AdaptLLM Cheng et al. (2024) is a LLaMA-1 model (7B parameters) pre-trained on 3.6M US court opinions from the FreeLaw project Gao et al. (2020). It converts raw legal texts into reading comprehension tasks, such as summarization, NLI, commonsense reasoning, and text completion, using regex-based patterns, and learns in a self-supervised manner.
SaulLM-7B SaulLM-7B Colombo et al. (2024) is a Mistral-based model (7B parameters) pre-trained on 94B tokens of legal text. It combines data from existing sources such as FreeLaw, the English MultiLegal Pile, EDGAR, and EuroParl, along with additional datasets including GovInfo, Law Stack Exchange, Australian legal corpora, EU and UK legislation, court transcripts, and USPTO filings. The corpus is curated from both public datasets and web-scraped legal resources to ensure broad legal coverage across jurisdictions.
@@ Appendix C Experimental Setup
We use all publicly available legal-specific pre-trained models from Hugging Face. To ensure fair comparison, we adopt the training configuration introduced by Chalkidis et al. (2022) for the LEDGAR and UNFAIR-ToS datasets: a learning rate of $3\text{e-}5$ for all nine encoder-base models and $1\text{e-}5$ for the encoder-large model PoL-BERT Henderson et al. (2022), consistent with the setting used for RoBERTa-large. All models are trained for up to 20 epochs with a batch size of 8, using early stopping with a patience of 3 based on development set performance. For UNFAIR-ToS, we use a maximum sequence length of 128, as in Chalkidis et al. (2022). However, we disable mixed-precision training (i.e., set fp16=False) to ensure stable training, which results in longer training times compared to Chalkidis et al. (2022). For LEDGAR, we reduce the maximum sequence length to 128 (from 512 used in Chalkidis et al. (2022)) to save computational resources and training time. This adjustment is necessary because, as discussed above, fp16 is disabled to ensure stable training, which leads to longer training times, and the LEDGAR dataset contains over 80k sentences, which is large. We observe only a marginal performance drop (0.1-0.4%), which we consider acceptable for efficiency. Each model is trained five times with different random seeds (1-5), and we report the test results of the best seed, following Chalkidis et al. (2022) for a fair baseline comparison. For the LEXDEMOD dataset, we follow the setup proposed by Sancheti et al. (2022), using a learning rate of $2\text{e-}5$ for all encoder-based legal-specific models, including PoL-BERT, consistent with their configuration for RoBERTa-large. We use a batch size of 8 and apply early stopping with a patience of 3. The maximum sequence length is set to 256, as in Sancheti et al. (2022). Each model is trained five times with different random seeds, and we report the average test performance across the three best seeds, following Sancheti et al. (2022) for a fair baseline comparison. For decoder-based generalist and legal-specific models, the parameters remain the same. However, for the LEDGAR dataset, the number of epochs for models such as Mistral, AdaptLLM, and SaulLM is reduced to 3 instead of 20 due to the long training time (22-24 hours per model). This adjustment is made because fp16 is disabled for stable training, and it is necessary for LEDGAR due to its large model size and the resource constraints. In addition to encoder-decoder based legal-specific model (LexT5), we adopt instruction-based fine-tuning, which aligns better with encoder-decoder models Wang et al. (2023). This approach pairs natural language prompts with clause inputs, enabling the model to generate the appropriate label(s) as output. The instruction templates used are listed in Figure 5. For UNFAIR-ToS, we directly use the prompt from Chalkidis (2023), originally designed for zero-shot prompting. For LEXDEMOD and LEDGAR, we design our own prompts inspired by that style. We evaluate model performance using micro-F1 ($\mu$-F1) and macro-F1 (m-F1) to account for class imbalance. Additionally, we report the arithmetic mean with standard deviation for micro-F1 ($\mu$-F1) and macro-F1 (m-F1) across tasks. All experiments are conducted on a single NVIDIA V100 GPU.
Appendix D Other generalist Models Considered for Exploratory Purposes
Here, we discuss the closed-source model (GPT-3.5-Turbo) results using prompting techniques, as reported by Chalkidis (2023). For UNFAIR-ToS, zero-shot prompting yields micro and macro F1-scores of 41.4 and 22.2, respectively. In few-shot prompting, performance improves, achieving micro and macro F1-scores of 64.7 and 32.5, though these scores remain lower than those achieved through legal-specific and generalist model fine-tuning techniques. A similar pattern occurs with LEDGAR, where zero-shot prompting results in micro and macro F1-scores of 70.1 and 56.7, respectively, while few-shot prompting produces scores of 62.1 and 51.1. Here, performance decreases with few-shot prompting, but task-specific fine-tuning using legal and generalist models remains the preferred approach. Thus, task-specific fine-tuning in the legal domain is essential for accurate legal contract classification.